# Model X-ray : Detect Backdoored Models via Decision Boundary

### v1

@@ 1 Introduction
Table 2: The backdoor attacks involved in our evaluations have covered diverse trigger patterns.
2 Related Work
2.1 Backdoor Attacks
$\begin{array}[]{r}\theta=\arg\min_{\theta}\mathbb{E}_{(x,y)\sim\mathcal{D}}\mathcal{L}(\hat{M}(x;\theta),y)\\ +\mathbb{E}_{\left(\hat{x},y_{t}\right)\sim\mathcal{\hat{D}}}\mathcal{L}(\hat{M}(\hat{x};\theta),y_{t}),\end{array}$ (1)
where $\mathcal{D}$ and $\mathcal{\hat{D}}$ denote the benign samples and trigger samples, respectively. $\mathcal{L}$ denotes the loss function, e.g., cross-entropy loss. The infected model functions normally on benign samples but yields a specific target prediction $y_{t}$ when presented with trigger samples $\hat{x}$. Backdoor attacks can be achieved by data poisoning and model modification, and we briefly introduce some related methods in the following part.
Data poisoning-based backdoor attacks primarily revolve around crafting trigger samples. Notably, BadNets (Gu et al., 2017) was a pioneering work that highlighted vulnerabilities of DNNs by employing visible squares as triggers. Afterward, various other visible trigger techniques have been explored: Blended (Chen et al., 2017) employs image blending to create trigger patterns, SIG (Barni et al., 2019) utilizes sinusoidal strips as triggers, and Low Frequency (LF) (Zeng et al., 2021) explores triggers in the frequency domain. Simultaneously, other research endeavors focus on achieving imperceptibility of the trigger patterns, including BPP (Wang et al., 2022) based on image quantization and dithering, WaNet (Nguyen & Tran, 2021) founded on image warping, and SSBA (Li et al., 2021a) achieved by image steganography. During the training stage, the attacker can leverage different poisoning ratios to balance the attack ability and performance degradation.
Apart from data poisoning-based attacks, there are some backdoor attacks that employ model modification techniques. TrojanNN (Liu et al., 2018) first proposes to optimize the trigger to ensure that the crucial neurons can attain their maximum values, LIRA (Doan et al., 2021) formulates malicious function as a non-convex, constrained optimization problem to learn invisible triggers through a two-stage stochastic optimization procedure, and Blind (Bagdasaryan & Shmatikov, 2021) modifies the training loss function to enable the model to learn the malicious function.
In this paper, we evaluate our method across both data poisoning-based attacks and model modification-based attacks, as shown in Table 2.
2.2 Backdoor Defenses
Comparably, post-training backdoor detection is model-level detection. Neural Cleanse (Wang et al., 2019) is the first post-training detection through anomaly analysis on the reversed trigger patterns. However, it requires access to the model’s inner information like parameters and gradients, which is also the limitation of other subsequent methods (Wang et al., 2019, 2020; Fu et al., 2023; Liu et al., 2019; Wang et al., 2023). Differently, detection work in black-box scenarios is extremely challenging (Chen et al., 2019; Xu et al., 2021; Guo et al., 2021; Dong et al., 2021), e.g. MNTD trains a meta-classifier based on features extracted from a large set of shadow models. However, its success heavily relies on the generalization capability of the attack settings from the shadow models to the actual backdoored models. Besides, it requires the soft label generated by the target model.
2.3 Decision Boundary of Deep Neural Networks
Most previous works depict decision boundaries by adversarial samples (He et al., 2018; Khoury & Hadfield-Menell, 2018) or sensitive samples (He et al., 2019). These methods are pivotal in identifying and understanding the contours of decision boundaries, as adversarial and sensitive samples are typically positioned along these critical junctures in the model’s decision-making process. However, obtaining these special samples requires access to the target model. Fortunately, Zhang et al. (Zhang et al., 2017) find that decision boundaries not only manifest near the data manifold but also within the convex hull created by pairs of data points. Leveraging this understanding, Somepalli et al. (Somepalli et al., 2022) introduce an innovative approach that utilizes only clean samples to map out the decision boundary to investigate reproducibility and double descent. Their method, which results in a two-dimensional (2D) map, offers an intuitive and accessible means of visualizing decision boundaries. In this paper, we utilize this technique to detect backdoored APIs in MLaaS scenarios.
3 Preliminaries
3.1 Recap of the Decision Boundary in (Somepalli et al., 2022)
After representing the space, we can calculate the bounds on the X-axis and the Y-axis, extended by a factor of $\eta$ in both the positive and negative directions along the corresponding axes, serving as a means to control the expansion range of the coordinate system. In the previous work (Somepalli et al., 2022), $\eta$ is set as 1 to investigate reproducibility, while we set $\eta$ as 5 to obtain a wider range of the decision boundary (see Fig. 3). Moreover, we can also determine density $S$ by constructing the set of points with a quantity of $S^{2}$ within the bounded range of the coordinate system using a grid generation method. Larger $S$ means higher resolution. With $S^{2}$ points, we can conduct the reverse process to get their tensor presentation, which can be fed to the API to fetch the corresponding prediction. We adopt different colors for different predictions to get the final 2D decision boundary.
Figure 3: Visual examples of the decision boundary used in (Somepalli et al., 2022) (left) and in this paper (right).
3.2 Threat Model
In this paper, the defender is a user who downloads or purchases the API from MLaaS clouds. As shown in Table 1, before usage, he aims to detect whether the target model is compromised by backdoor attacks including data poisoning-based attacks and model modification-based attacks. Besides, it is better to determine the target attacked label. In terms of the defender’s capabilities, there is no access to the training data and inference data. Notably, the only capacity of the defender is the hard prediction generated by the target API.
4 Method
In this section, we first provide an intriguing observation on the decision boundary of clean models and backdoor models. Based on the observation, we designate two strategies for backdoor detection via the decision boundary. Finally, we showcase that we can determine the target attacked label, if only one single label is infected.
Figure 4: Visual examples of decision boundaries of the clean model and different backdoored models on CIFAR-10 and ImageNet-10.
4.1 An Intriguing Observation
As shown in Fig. 4, we provide the decision boundary of the clean model and different backdoored models (infected by BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), LF (Zeng et al., 2021), BPP (Wang et al., 2022), TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan & Shmatikov, 2021)) on CIFAR-10 and ImageNet-10 dataset. We observe that the prediction distribution within the decision boundary of the backdoored model is more gathered, and triple samples are encircled by a large area of the target label. We explain it may be the shortcut effect caused by backdoor attacks. The shortcuts leading to the attack target label in the backdoor model has been confirmed in previous research, that is, through optimization methods, smaller perturbations can be found to cause other labels to be misclassified as target labels (Wang et al., 2019). Afterward, Rajabi et al. (Rajabi et al., 2023) quantifies this effect by introducing the concept of a certified radius (Cohen et al., 2019), which estimates the distance to a decision boundary by perturbing samples with Gaussian noise with a predetermined mean and variance. Notably, trigger samples are observed to be relatively farther from a decision boundary compared to clean samples, which can support why the large region is dominated by injected prediction.
Figure 5: Illustration on calculation of RE and ATS.
4.2 Two Strategies for Backdoor Detection via the Decision Boundary
As discussed above, in contrast to clean models, backdoor models have anomalous decision boundaries. Therefore, backdoor detection can be transformed into anomaly detection on the decision boundary. To achieve this, we propose two strategies for backdoor detection via the decision boundary, namely, based on Rényi Entropy (RE) and Areas dominated by Triple Samples (ATS), respectively. In the following part, we will introduce the two strategies in detail, which we hope sheds some light on anomaly detection. Notably, other strategies are also applicable.
4.2.1 Backdoor Detection based on Rényi Entropy
With the technique mentioned above, we can plot $N$ decision boundaries $\mathbf{B}=\{\mathcal{B}_{1},...,\mathcal{B}_{k},...,\mathcal{B}_{N}\}$, where $\mathcal{B}_{k}$ is plotted along the plane spanned by triple samples $T_{k}=(x_{1},x_{2},x_{3})_{k}$. Specifically, let $S_{k}=\{x_{ij}|(i,j)\in\mathcal{B}_{k}\}$ be the set of points in the $\mathcal{B}_{i}$, where $(i,j)$ is the coordinations of $x$ in $\mathcal{B}_{k}$. Then, we feed $S_{k}=\{x_{ij}|(i,j)\in\mathcal{B}_{k}\}$ to the target model API $M$ to obtain the corresponding hard labels $L_{k}=\{l_{ij}|(i,j)\in\mathcal{B}_{k}\}$, which are further used to obtain the final colorful decision boundary $\mathcal{B}_{k}$ for evaluation.
$\small p_{m}=\frac{A(l_{m})}{A(\mathcal{B}_{k})},$ (2)
where $l_{m}$ denotes the $m$-th class label in the dataset. $A(l_{m})$ and $A(\mathcal{B}_{k})$ denote the areas of $m$-th class and the areas of entire decision regions, respectively. In Fig. 5 (left), for example, $p_{3}={(A(\textbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\textcircled{3}}})+A(\textcircled{3}))}/{A(\mathcal{B}_{k})}$. To indirectly evaluate the gathering degree of the decision boundary, we calculate Rényi Entropy (RE) of label probability distribution $\mathcal{P}_{k}$:
$\small RE(\mathcal{P}_{k})=H_{\alpha}(\mathcal{P}_{k})=\frac{1}{1-\alpha}\log\left(\sum_{m=1}^{n}p_{m}^{\alpha}\right),$ (3)
4.2.2 Backdoor Detection based on Areas dominated by Triple Samples
$\small ATS(\mathcal{B}_{k})={\frac{A(T_{k})}{A(\mathcal{B}_{k})}}={\frac{\sum_{x\in{(x_{1},x_{2},x_{3})}}{A(x)}}{A(\mathcal{B}_{k})}},$ (4)
where ${A(T_{k})}$ denotes the total areas dominated by triple samples. As shown in the left of Fig. 5, $ATS(\mathcal{B}_{k})=(A(\textbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\textcircled{1}}})+A(\textbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\textcircled{2}}})+A(\textbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\textcircled{3}}}))/A(\mathcal{B}_{k})$. However, we find there are some special cases. As shown in Fig. 5 (right), one of the triple samples belongs to the target attacked label, causing an abnormally large $A(\textbf{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\textcircled{1}}})$⃝ ). In practice, we cannot determine whether the labels of triple samples are injected. For this, we append an additional constraint for ATS, namely, ${A(x)}<{A(\mathcal{B}_{k})}\cdot t$, where $t=0.5$ by default. Based on ATS, we propose a detection strategy called Ours-ATS. Intuitively, the large ATS means robust classification on the clean images, and vice versa.
Figure 6: The label probability distribution within decision boundaries of clean and backdoor models on CIFAR-10 and ImageNet-10, both of whose infected labels are 0.
4.3 Determine the Target Label
@@
Table 3: The performance of Model X-ray across different attacks, datasets, and architectures. The last two columns show the worst and the average performance among different attacks. The best results are in bold.
5 Experiment
5.1 Experimental Settings
Datasets and Architectures. The datasets include CIFAR-10 (Krizhevsky et al., 2009), CIFAR-100 (Krizhevsky et al., 2009), GTSRB (Houben et al., 2013), and ImageNet-10 (Howard, ), a subset of ten classes from ImageNet (Deng et al., 2009). Besides, we employ four different architectures: PreActResNet-18 (He et al., 2016), MobileNet-V3-Large (Howard et al., 2019), PreActResNet-34 (He et al., 2016), and ViT-B-16 (Dosovitskiy et al., 2020). These architectures encompass both Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs) and span across various network sizes, including small, medium, and large networks.
Implementation Details. For the model to be evaluated, we plot decision boundaries by random samples triplet with expansion factor $\eta=5$ and density $S=100$, number of plots $N=20$. For the attack baselines, we evaluate our method against seven backdoor attacks, including BadNets (Gu et al., 2017), SSBA (Li et al., 2021a), LF (Zeng et al., 2021), BPP (Wang et al., 2022), TrojanNN (Liu et al., 2018), LIRA (Doan et al., 2021), and Blind (Bagdasaryan & Shmatikov, 2021). We follow an open-sourced backdoor benchmark BackdoorBench (Wu et al., 2022) for the training settings of these attacks and conduct all-to-one attacks by default. As shown in Table 2, the attacks in our experiments contain diverse complex trigger pattern types. In this paper, we focus on backdoor detection for MLaaS. We compare with three post-training detection methods, namely, Neural Cleanse (Wang et al., 2019), MNTD (Xu et al., 2021) and MM-BD (Wang et al., 2023).
@@ 5.1 Experimental Settings
Figure 7: The average RE ($\alpha=10$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10 datasets. We observe that backdoor models have significantly smaller RE than clean models.
@@
Figure 8: The average ATS ($t=0.5$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10 datasets. We observe that backdoor models have significantly smaller ATS than clean models.
5.2 The Effectiveness of Model X-ray
As shown in Table 3, in most cases, Model X-ray outperforms the baseline methods across different backdoor attacks, datasets, and architectures. Notably, the training results for the meta-classifier exhibit some degree of instability. The reported MNTD results represent the best outcomes selected from five independent runs. Neural Cleanse demonstrates promising performance in the majority of scenarios. However, occasional failures may arise when it incorrectly identifies a trigger for a clean model, leading to convergence in local optima. Crucially, Model X-ray stands out as the sole effective method for MLaaS scenarios, whereas Neural Cleanse necessitates access to the model’s parameters, and MNTD relies on soft predictions from the target API. In Fig. 7 and Fig. 8, we provide visual illustrations of the average RE and ATS among both clean and backdoored models. In most cases, a clear distinction is evident, except for LF, BBP, and TrojanNN on CIFAR-10, which aligns with the quantitative results presented in Table 3. Employing improved analysis strategies may help address this issue.
5.3 The Efficiency of Model X-ray
In Table 4, we show the number of benign samples that the defender needs. Both Neural Cleanse and MNTD necessitate a certain proportion of benign data (e.g., 5% of the benign dataset) to complement their defense mechanisms. In contrast, our method necessitates only three benign samples to plot a decision boundary, and with $N$ set to 20, only 60 clean samples are required, which is already sufficient to ensure the effectiveness of our detection. In addition, we compare the average evaluation execution time of each method on CIFAR-10 dataset in Table 5. The experiment is conducted on one NVIDIA RTX A6000. Specifically, Neural Cleanse requires a trigger reverse engineering optimization process, MM-BD also requires an optimization process and MNTD requires preparation that generates a large set of shadow models (1024 clean models and 1024 attack models) to train a meta-classifier. In contrast, our method does not demand any optimization or training processes, making it a versatile and plug-and-play solution.
Method CIFAR-10 GTSRB CIFAR-100 ImageNet-10
Neural Cleanse 2500 1332 2500 473
MNTD 2500 1332 2500 473
MM-BD 0 0 0 0
Ours 60 60 60 60
Table 4: Benign samples required for different methods.
Method Neural Cleanse MNTD $\dagger$ MNTD MM-BD Ours
Time (s) 243.4 44268.6 0.08 75.2 36.5
Table 5: Execution time on CIFAR-10. $\dagger$ means the training time.
5.4 threshold
Since Model X-ray maps the model to a linearly separable space and defenders make judgments based on a threshold $\gamma$. Observing the RE and ATS of clean and backdoor models with different architectures on the same dataset (see Fig. 9 and Fig. 10), we find that detection between models in the same dataset on different architectures exhibits transferability. This allows us to empirically determine a threshold $\gamma$.
(a) PreActResNet18
(b) MobileNet-L
(c) PreActResNet34
(d) VGG19
Figure 9: The average RE($\alpha=10$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10 on different architectures.
(a) PreActResNet18
(b) MobileNet-L
(c) PreActResNet34
(d) VGG19
Figure 10: The average ATS($t=0.5$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10 on different architectures.
5.5 Ablation Study
The Influence of the Hyper-parameters. $N$ is the number of decision boundary plots and $S$ is the density of decision boundaries, which are critical to the evaluation efficiency. Here, we investigate Model X-ray ’s performance under fixed $N=20$ with $S$ ranging from 60 to 140 and under fixed $S=100$ with $N$ ranging from 5 to 40. Fig. 11 shows that lower $N$ and $S$ will slightly degrade the performance of Model X-ray on CIFAR-10, which is still acceptable.
Figure 11: The influence of the number of plots $N$ and density $S$.
Besides, we investigate the impact of parameters in two indicators, i.e., $\alpha$ in RE and $t$ in ATS. As shown in Fig. 12, different $\alpha$ has a neglectable effect on Ours-RE, while $t$ larger than 0.5 is better for Ours-ATS.
Figure 12: The influence of the parameters $\alpha$ and $t$.
Figure 13: The influence of the poisoning ratio.
The Impact of Different Attack Strategies. Besides the default all-to-one attack strategy, we consider the X2X attack strategy (Xiang et al., 2023), with arbitrary numbers of source classes each assigned with an arbitrary target class, including X-to-X attack, X-to-one attack, and one-to-one attack. We adopt different attack strategies to conduct BadNets on CIFAR-10. For each strategy, we train 10 models for evaluation. Table 6 shows that Model X-ray remains effective under different attack strategies, especially based on ATS (i.e., Ours-ATS). Furthermore, we provide some visual examples of the corresponding decision boundary in Fig. 14. Taking 10-to-10 as an example, wherein all labels are injected, the area for triple clean samples shrinks, indicating the degraded robustness of clean samples.
@@ 5.5 Ablation Study
Table 6: The performance under different attack strategies.
Figure 14: Decision boundaries under different attack strategies.

### v2

@@ 5.1. Experimental Settings
Table 2. The performance of Model X-ray across different attacks, datasets, and architectures. The last two columns show the worst and the average performance among different attacks. The best results are in bold.
@@
Figure 7. The average RE ($\alpha=10$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10 datasets. We observe that backdoor models have significantly smaller RE than clean models.
@@
Figure 8. The average ATS ($t=0.5$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10, CIFAR-100, GTSRB, and ImageNet-10 datasets. We observe that backdoor models have significantly smaller ATS than clean models.
5.2. The Effectiveness of Model X-ray
As shown in Table 2, in most cases, Model X-ray outperforms the baseline methods across different backdoor attacks, datasets, and architectures. MNTD is difficult to generalize attack settings from the shadow models to the actual backdoored models. Neural Cleanse performs well in the majority of scenarios. However, occasional failures may arise when it incorrectly identifies a trigger for a clean model, leading to convergence in local optima. MM-BD demonstrates promising performance on small-scale architectures, but its performance drops significantly on larger architectures. In Fig. 7 and Fig. 8, we present visual illustrations of the average RE and ATS values for both clean and backdoored models. In most cases, a clear distinction is evident between clean and backdoored models. The ROC curves of Ours-RE and Ours-ATS can be found in the supplementary material.
Table 3. The performance under different attack strategies.
@@ 5.2. The Effectiveness of Model X-ray
(a) PreActResNet18
(b) MobileNet-L
(c) PreActResNet34
(d) VGG19
Figure 9. The average RE ($\alpha=10$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10 on different architectures.
(a) PreActResNet18
(b) MobileNet-L
(c) PreActResNet34
(d) VGG19
Figure 10. The average ATS ($t=0.5$) for clean and backdoor models injected by seven backdoor attacks in CIFAR-10 on different architectures.
Figure 11. Decision boundaries under different attack strategies.
Besides the default all-to-one attack strategy, we consider attack strategies (Xiang et al., 2023) with arbitrary numbers of source classes each assigned with an arbitrary attack target class, including X-to-X attack, X-to-one attack, and one-to-one attack. We adopt different attack strategies to conduct BadNets on CIFAR-10. For each strategy, we train 10 models for evaluation. Table 3 shows that Model X-ray remains effective under different attack strategies, especially based on ATS (i.e., Ours-ATS). Although multi-target attacks lower the performance of the proposed method, we outperform the baseline methods by a large margin in most cases. Furthermore, we provide some visual examples of the corresponding decision boundary in Fig. 11. In X-to-one and one-to-one attacks, where the attack target is a single class, both Ours-RE and Ours-ATS achieve precise detection and identification of the target class. In X-to-X attack, where there are multiple classes for both source and attack targets, the performance of Ours-RE declines with an increasing number of attack target classes, which is acceptable. The computation of Ours-RE relies on the entropy of class labels, where it can still detect the presence of multiple attack target classes in the decision boundary, despite the performance drop. Furthermore, areas dominated by triple clean samples shrink, which explains why Ours-ATS achieves good performance in such scenarios.
5.3. Evaluations on Open-source Benchmarks
To mitigate the impact of incidental factors in our training, we also evaluated our method on the backdoored models pre-trained on an open-source benchmark (Wu et al., 2022). Specifically, we perform detection on pre-trained backdoored models injected with seven backdoor attacks across CIFAR-10, GTSRB, and CIFAR-100 datasets using the PreActResNet-18 architecture, which can be downloaded from Open-source benchmarks (Bac, 2023).
Given a target model $C_{\theta}$, Model X-ray map the model $C_{\theta}$ to a linearly separable space, defenders can make judgments through average RE and ATS based on a threshold $\gamma$ :
[…]
As shown in Fig. 9 and Fig. 10, for the same dataset (taking CIFAR-10 as an example), we find that the relationship of RE and ATS between clean and backdoor models exhibits consistency. This allows us to determine an estimated threshold $\bar{\gamma}$ based on a small set of models:
(6) $\bar{\gamma}=\frac{1}{N}\sum_{m=1}^{N}\arg\max_{\gamma\in\Gamma}\frac{2\times\left(\text{ precision }_{\gamma}\times\text{ recall }_{\gamma}\right)}{\left(\text{ precision }_{\gamma}+\operatorname{recall}_{\gamma}\right)}.$
Based on thresholds $\bar{\gamma}$ (e.g., for Ours-RE CIFAR-10: $0.873$, GTSRB: $2.040$, CIFAR-100: $1.194$; for Ours-ATS, CIFAR-10: $0.184$, GTSRB: $0.134$, CIFAR-100: $0.040$), the detection accuracy on CIFAR-10 is 87.5%, on GTSRB is 93.75% and on CIFAR-100 is 100%. The visualized decision boundaries can be found in the supplementary material.
5.4. The Efficiency of Model X-ray
Neural Cleanse and MM-BD necessitate access to the model’s parameters, and MNTD relies on logit outputs from the target model. Model X-ray detects the backdoored model solely by predicted hard labels of clean inputs from the model. In Table 4, we show the number of benign samples that the defender needs. Both Neural Cleanse and MNTD necessitate a certain proportion of benign data (e.g., 5% of the benign dataset) to complement their defense mechanisms, MM-BD does not require any clean data. Our method necessitates only three benign samples to plot a decision boundary, and with $N$ set to 20, only 60 clean samples are required, which is already sufficient to ensure the effectiveness of our detection.
In addition, we compare the average inference time of each method in Table 5. The experiment is conducted on one NVIDIA RTX A6000. Specifically, Neural Cleanse requires a trigger reverse engineering optimization process for each class, MM-BD also requires a margin statistical process to obtain a maximum margin statistic for each class, and MNTD requires preparation that generates a large set of shadow models (1024 clean models and 1024 attack models) to train a meta-classifier. In contrast, our method eliminates the need for any optimization or training processes, making it a versatile plug-and-play solution that functions as a lightweight diagnostic scanning tool.
Table 4. Benign samples required for different methods.
Method CIFAR-10 GTSRB CIFAR-100 ImageNet-10
Neural Cleanse 2500 1332 2500 473
MNTD 2500 1332 2500 473
MM-BD 0 0 0 0
Ours 60 60 60 60
Table 5. The average inference time(sec) for different methods. $\dagger$ means the training time(sec).
Method Neural Cleanse MNTD $\dagger$ MNTD MM-BD Ours
CIFAR-10 243.4 44268.6 0.06 75.2 36.5
GTSRB 628.5 53409.0 0.05 334.8 34.6
CIFAR-100 2431.7 46680.9 0.06 829.5 36.0
ImageNet-10 1471.0 73632.2 1.5 414.4 112.3
Figure 12. The influence of the number of plots $N$ and point density $S$.
Figure 13. The influence of the parameters $\alpha$ and $t$.
Figure 14. The influence of the poisoning ratio.
5.5. Ablation Study
The Influence of the Hyper-parameters. $N$ is the number of decision boundary plots and $S$ is the density of decision boundaries, which are critical to the evaluation efficiency. Here, we investigate Model X-ray’s performance under fixed $N=20$ with $S$ ranging from 60 to 140 and under fixed $S=100$ with $N$ ranging from 5 to 40. Fig. 12 shows that lower $N$ and $S$ will slightly degrade the performance of Model X-ray on CIFAR-10, which is still acceptable. Besides, we investigate the impact of parameters in two indicators, i.e., $\alpha$ in RE and $t$ in ATS. As shown in Fig. 13, different $\alpha$ has a neglectable effect on Ours-RE, while $t$ larger than 0.5 is better for Ours-ATS.